# Stock Market Prediction Using Node Transformer Architecture Integrated with BERT Sentiment Analysis

### 06 Mar 2026 v1

@@ Abstract
Stock market prediction presents considerable challenges for investors, financial institutions, and policymakers operating in complex market environments characterized by noise, non-stationarity, and behavioral dynamics. Traditional forecasting methods, including fundamental analysis and technical indicators, often fail to capture the intricate patterns and cross-sectional dependencies inherent in financial markets. This paper presents an integrated framework combining a node transformer architecture with Bidirectional Encoder Representations from Transformers (BERT) based sentiment analysis for stock price forecasting. The proposed model represents the stock market as a graph structure where individual stocks form nodes and edges capture relationships including sectoral affiliations, correlated price movements, and supply chain connections. A fine-tuned BERT model extracts sentiment information from social media posts and combines it with quantitative market features through attention-based fusion mechanisms. The node transformer processes historical market data while capturing both temporal evolution and cross-sectional dependencies among stocks. Experiments conducted on 20 S&P 500 stocks spanning January 1982 to March 2025 demonstrate that the integrated model achieves a mean absolute percentage error (MAPE) of 0.80% for one-day-ahead predictions, compared to 1.20% for ARIMA and 1.00% for LSTM. The inclusion of sentiment analysis reduces prediction error by 10% overall and 25% during earnings announcements, while the graph-based architecture contributes an additional 15% improvement by capturing inter-stock dependencies. Directional accuracy reaches 65% for one-day forecasts. Statistical validation through paired t-tests confirms the significance of these improvements ($p<0.05$ for all comparisons). The model demonstrates robustness during high-volatility periods, maintaining MAPE below 1.5% where baseline models exceed 2%.
@@ II-D Graph Neural Networks
Graph Neural Networks (GNNs) capture interdependencies by modeling markets as graphs where nodes represent companies and edges represent relationships such as sectoral ties, price correlations, or supply chain connections.
Chen et al. [22] proposed a graph convolutional feature-based CNN model combining graph convolutional networks with dual CNNs to capture market-level and stock-level features simultaneously. Testing on Chinese stocks demonstrated superior performance compared to non-graph approaches. Wang et al. [23] introduced a multi-graph convolutional neural network defining both static and dynamic correlation graphs. Testing on 42 Chinese market indices showed average prediction error reduction of 5.11% compared to LSTM and attention-based baselines.
@@ II-F Hybrid Models and Research Gaps
Despite progress along each of these fronts, critical gaps persist at their intersections. Few approaches combine graph-based inter-stock modeling with transformer temporal processing. Integration of unstructured text with structured numerical data remains limited. Performance often deteriorates during volatile conditions, with accuracy varying by 40% between stable and turbulent periods [11]. Computational demands present barriers to real-time deployment.
These gaps motivate the present work: an integrated node transformer with BERT-based sentiment analysis, capturing both stock interconnections and long-range dependencies while incorporating sentiment as a predictive signal.
@@ III-A Financial Market Dataset
The selected companies span multiple sectors: technology (Apple, Microsoft, Salesforce, Netflix), financial services (JPMorgan Chase, Visa), healthcare (Johnson & Johnson, UnitedHealth Group, Pfizer), retail (Walmart, Home Depot, McDonald’s), energy (ExxonMobil, Chevron), consumer goods (Procter & Gamble, Coca-Cola, Nike), telecommunications (Verizon), and industrials (Boeing, Caterpillar). For companies with initial public offerings after 1982, data collection begins at the first available trading date. For example, Salesforce data begins in June 2004 following its IPO. The effective analysis period for each stock corresponds to its available trading history within the 1982-2025 window. This cross-sectional design ensures the dataset reflects behavior across multiple industries and economic cycles, though the selection of current index constituents introduces survivorship bias that is acknowledged in Section VI.
@@ III-C Dataset Partitioning
Temporal splits were applied to prevent look-ahead bias. The training set comprises January 1, 1982 to December 31, 2010 (70% of temporal range). The validation set spans January 1, 2011 to December 31, 2016 (15%). The test set covers January 1, 2017 to March 31, 2025 (15%).
For sentiment data, partitioning aligns with financial splits for the overlapping period. Since X (formerly Twitter) data begins in 2007, the sentiment training period covers January 2007 to December 2010. For pre-2007 financial data, sentiment features are set to neutral (zero). The sentiment integration component is primarily evaluated on the test set (2017-2025) where sentiment data is abundant.
@@ IV-C2 Temporal Encoding
where $t$ is the trading day index, $k\in\{0,\ldots,d/2-1\}$ is the dimension index, and $d=512$ is the model dimension. Lower dimensions capture high-frequency patterns; higher dimensions capture low-frequency patterns.
@@ IV-D2 Fine-Tuning Process
where $\alpha_{t}$ is class weight (inverse frequency), $p_{t}$ is predicted probability for correct class, and $\gamma=2$ focuses training on hard examples by down-weighting easily classified instances. Fig. 4 illustrates the complete sentiment extraction pipeline from raw input to final score.
@@ IV-G Training Process
Throughout all three stages, optimization uses Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The learning rate schedule begins with a linear warmup over 4000 steps followed by cosine decay to zero. A batch size of 32 with gradient accumulation over 4 steps yields an effective batch size of 128, balancing memory constraints with gradient stability. Table III summarizes the key hyperparameters for both the node transformer and BERT components.
@@ V-B Evaluation Metrics
The prediction target $y_{i,t+h}$ is the closing price $C_{t+h}$ at horizon $h\in\{1,5,20\}$ trading days. Price-level prediction provides intuitive interpretation of forecast accuracy and aligns with practical trading applications. Directional accuracy is derived from the implied price movement relative to the current price. Five complementary metrics are used to assess model performance from different angles, covering magnitude accuracy, scale sensitivity, directional correctness, relative forecasting skill, and cross-sectional ranking ability.
@@ V-C Baseline Models
Eight baseline models spanning statistical, classical machine learning, and deep learning paradigms were trained on identical data splits to ensure fair comparison. All hyperparameters were tuned via grid search on the validation set (2011-2016), with final performance reported on the held-out test set (2017-2025).
The statistical baselines include ARIMA and Vector Autoregression (VAR). ARIMA orders $(p,d,q)$ were selected via AIC with $p,q\in[0,5]$ and $d\in[0,2]$, fitting a separate model per stock. VAR extends this to the multivariate setting, modeling cross-stock dependencies with lag order selected via the Bayesian Information Criterion (BIC) up to 10 lags.
Three classical machine learning models were evaluated. Random Forest used 100 trees with maximum depth 10 and the same 17-dimensional feature vector as the proposed model. Support Vector Regression (SVR) employed a Radial Basis Function (RBF) kernel with regularization parameter $C$ and kernel width $\gamma$ selected via grid search. XGBoost used 1000 estimators with early stopping on validation loss to prevent overfitting, and its gradient boosting framework provides a strong tree-based benchmark.
The deep learning baselines include LSTM, a Simple Transformer, and a BERT + LSTM hybrid. The LSTM used 2 layers with 128 hidden units and a 60-day lookback window, representing the most widely adopted recurrent architecture for financial time series. The Simple Transformer used 4 layers, 8 attention heads, and 256-dimensional embeddings, providing a direct comparison that isolates the contribution of graph structure and sentiment integration in the proposed model. The BERT + LSTM hybrid combined BERT-extracted sentiment features with an LSTM price model through simple concatenation, serving as an ablation baseline that tests whether the attention-based fusion mechanism in the proposed architecture provides meaningful improvement over naive feature combination.

### 03 Apr 2026 v2

@@ Abstract
Stock market prediction presents considerable challenges for investors, financial institutions, and policymakers operating in complex market environments characterized by noise, non-stationarity, and behavioral dynamics. Traditional forecasting methods, including fundamental analysis and technical indicators, often fail to capture the intricate patterns and cross-sectional dependencies inherent in financial markets. This paper presents an integrated framework combining a node transformer architecture with Bidirectional Encoder Representations from Transformers (BERT) based sentiment analysis for stock price forecasting. The proposed model represents the stock market as a graph structure where individual stocks form nodes and edges capture relationships including sectoral affiliations, correlated price movements, and supply chain connections. A fine-tuned BERT model extracts sentiment information from social media posts and combines it with quantitative market features through attention-based fusion mechanisms. The node transformer processes historical market data while capturing both temporal evolution and cross-sectional dependencies among stocks. Experiments conducted on 20 S&P 500 stocks spanning January 1982 to March 2025 demonstrate that the integrated model achieves a mean absolute percentage error (MAPE) of 0.80% for one-day-ahead predictions, compared to 1.20% for ARIMA and 1.00% for LSTM. The inclusion of sentiment analysis reduces prediction error by 10% overall and 25% during earnings announcements, while the graph-based architecture contributes an additional 15% improvement by capturing inter-stock dependencies. Directional accuracy reaches 65% for one-day forecasts. Statistical validation through paired t-tests confirms the significance of these improvements ($p<0.05$ for all comparisons). The model maintains lower error during high-volatility periods, achieving MAPE of 1.50% while baseline models range from 1.60% to 2.10%.
@@ II-D Graph Neural Networks
Graph Neural Networks (GNNs) capture interdependencies by modeling markets as graphs where nodes represent companies and edges represent relationships such as sectoral ties, price correlations, or supply chain connections. Kipf and Welling [29] introduced the graph convolutional network (GCN) framework, which aggregates features from neighboring nodes through symmetric normalization of the adjacency matrix, providing a scalable approach to semi-supervised learning on graph-structured data. This foundational architecture enabled subsequent applications to financial markets, where inter-stock relationships form natural graph topologies.
Chen et al. [22] proposed a graph convolutional feature-based CNN model combining graph convolutional networks with dual CNNs to capture market-level and stock-level features simultaneously. Testing on Chinese stocks demonstrated superior performance compared to non-graph approaches. Wang et al. [23] introduced a multi-graph convolutional neural network defining both static and dynamic correlation graphs. Testing on 42 Chinese market indices showed average prediction error reduction of 5.11% compared to LSTM and attention-based baselines. Beyond equity markets, graph-based forecasting methods have been applied to other financial domains. Le et al. [31] developed an entropy-based GCN framework for real estate market prediction that segments spatial regions into anomalous and normal clusters before applying separate forecasting models, reducing prediction error across multiple horizons on U.S., U.K., and Australian housing data. A related study by Le et al. [32] combined entropy-based multivariate clustering with sentiment extracted from news headlines and macroeconomic indicators, demonstrating that this segmentation approach improves forecasting accuracy when paired with deep learning architectures such as transformers and LSTMs. These cross-domain results underscore the generality of graph-based representations for capturing complex dependencies in financial data.
@@ II-F Hybrid Models and Research Gaps
More recent work has explored integrating transformer attention with sentiment signals and privacy-preserving training strategies. Nejatbakhsh and Aliasgari [30] combined LSTM networks with transformer-based attention and FinBERT-derived sentiment in a federated learning framework, achieving an average $R^{2}$ of 0.91 across ten technology stocks. Their approach demonstrated that sentiment integration improves short-term forecasting accuracy, though the architecture treats each stock independently without modeling inter-stock relationships, and the evaluation was restricted to a single sector. These findings reinforce the value of multimodal fusion while highlighting the continued need for architectures that simultaneously capture cross-asset dependencies.
Despite progress along each of these fronts, critical gaps persist at their intersections. Few approaches combine graph-based inter-stock modeling with transformer temporal processing within a unified architecture. While sentiment integration has been explored alongside recurrent and attention-based models, simultaneous learning of inter-stock graph structure, long-range temporal patterns, and textual sentiment remains largely unaddressed. Performance often deteriorates during volatile conditions, with accuracy varying by 40% between stable and turbulent periods [11]. Computational demands present additional barriers to real-time deployment.
The present work addresses these gaps through integration rather than through the introduction of a fundamentally new learning primitive. Each individual component, namely graph structure learning, transformer-based attention, and BERT sentiment analysis, draws on established techniques. The contribution lies in the architectural design required to combine them within a single end-to-end framework. The node transformer must compute attention simultaneously across temporal positions and graph neighborhoods, which requires adapting the standard self-attention mechanism to operate over a three-dimensional input tensor rather than a flat sequence. The adaptive gating mechanism that fuses continuous sentiment scores with numerical price features handles temporal sparsity and missing data without degrading the price-based pathway, a challenge that simple feature concatenation does not address. The jointly learned edge weights provide an interpretable structural layer connecting model internals to domain knowledge about market relationships, offering a qualitative advantage that pure performance comparisons cannot capture. This integrative design philosophy follows a productive tradition in deep learning, where combining known components in carefully engineered architectures has yielded substantial advances over any single technique deployed in isolation.
@@ III-A Financial Market Dataset
The selected companies span multiple sectors: technology (Apple, Microsoft, Salesforce, Netflix), financial services (JPMorgan Chase, Visa), healthcare (Johnson & Johnson, UnitedHealth Group, Pfizer), retail (Walmart, Home Depot, McDonald’s), energy (ExxonMobil, Chevron), consumer goods (Procter & Gamble, Coca-Cola, Nike), telecommunications (Verizon), and industrials (Boeing, Caterpillar). For companies with initial public offerings after 1982, data collection begins at the first available trading date. For example, Salesforce data begins in June 2004 following its IPO. The effective analysis period for each stock corresponds to its available trading history within the 1982-2025 window. This cross-sectional design ensures the dataset reflects behavior across multiple industries and economic cycles. The 20-stock universe was chosen deliberately to balance temporal depth with cross-sectional coverage: a smaller, well-established set of equities allows for training on over four decades of data with complete and reliable records, which would be difficult to achieve with a much larger or historically reconstructed universe where data quality varies substantially. This design is best understood as a proof-of-concept demonstrating the architecture’s capacity to learn inter-stock dependencies and integrate sentiment signals, rather than a comprehensive market simulation. The selection of current index constituents does introduce survivorship bias, which is examined in detail in Section VI.
@@ III-C Dataset Partitioning
Temporal splits were applied to prevent look-ahead bias, following the standard practice in financial time series forecasting. Because stock prices exhibit serial dependence and evolving distributional properties, random or stratified partitioning schemes such as k-fold cross-validation would allow future observations to inform predictions of earlier time steps, producing artificially inflated accuracy estimates that do not reflect realistic forecasting conditions. In a forecasting context, the model must never train on data that postdates the prediction target; violating this constraint renders performance metrics unreliable regardless of their magnitude. Chronological partitioning ensures that the model is evaluated exclusively on data that postdates the training window, mirroring the information constraints faced by a real-time forecaster.
The training set comprises January 1, 1982 to December 31, 2010 (70% of temporal range). The validation set spans January 1, 2011 to December 31, 2016 (15%), serving as both a hyperparameter tuning period and a temporal buffer that separates training from test data by six calendar years. The test set covers January 1, 2017 to March 31, 2025 (15%), deliberately spanning multiple distinct market regimes including the pre-pandemic bull market (2017–2019), the COVID-19 crash and recovery (2020–2021), and the inflationary tightening period (2022–2024). The inclusion of the COVID-19 pandemic in the test set rather than in training provides a rigorous assessment of the model’s ability to generalize to market conditions qualitatively different from those observed during training. The volatility regime analysis in Section V-E confirms that the model maintains reasonable accuracy across these diverse conditions, with MAPE rising from 0.70% in low-volatility periods to 1.50% in high-volatility periods, a degradation consistent with the inherent difficulty of forecasting during market stress rather than a failure of generalization.
The full 1982-2025 training window is retained because the price-based components of the model, including the node transformer, graph structure learning, and temporal encoding, benefit substantially from exposure to diverse market regimes spanning over four decades. Restricting training to the post-2007 period would sacrifice the 1982-2006 data that contains the 1987 crash, the dot-com bubble, and the 2008 financial crisis, all of which are critical for learning robust representations of market dynamics.
For sentiment data, partitioning aligns with financial splits for the overlapping period. Since X (formerly Twitter) data begins in 2007, the sentiment training period covers January 2007 to December 2010. For pre-2007 financial data, sentiment features are set to neutral (zero). This zero-padding is handled by the adaptive gating mechanism (Section IV-E), which learns to down-weight the sentiment channel when its signal is absent, effectively allowing the model to operate in a price-only mode for the pre-2007 training window and transition to multimodal operation when sentiment data becomes available. The sentiment integration component is primarily evaluated on the test set (2017-2025), where sentiment data is abundant and the model can fully leverage both data streams.
@@ IV-B Graph Representation of the Stock Market
At each training step, the model receives a batch of temporal windows. For a single sample, the input is a three-dimensional tensor $\mathbf{X}\in\mathbb{R}^{T\times N\times d_{\text{in}}}$, where $T=252$ is the sequence length (trading days), $N=20$ is the number of stock nodes, and $d_{\text{in}}=17$ is the per-stock feature dimension after concatenating raw OHLCV data with technical indicators. The edge weight matrix $\mathbf{E}\in\mathbb{R}^{N\times N}$ is initialized once from Equation 12 and shared across all time steps within a given forward pass. At each time step $t$, the node transformer processes the slice $\mathbf{X}_{t}\in\mathbb{R}^{N\times d_{\text{in}}}$ jointly across all $N$ nodes, with the attention mechanism modulated by $\mathbf{E}$ so that information flows preferentially along stronger graph connections. The edge weights are then updated via Equation 13 at each transformer layer, allowing the graph structure to co-evolve with the learned node representations as the input propagates through the network.
@@ IV-C2 Temporal Encoding
where $t$ is the trading day index, $k\in\{0,\ldots,d/2-1\}$ is the dimension index, and $d=512$ is the model dimension.
A clarification on the role of this encoding is warranted, as the use of sinusoidal functions may appear counterintuitive for financial data that does not follow periodic patterns. The temporal encoding does not model or approximate the shape of the price series. Instead, it provides the attention mechanism with information about where each observation falls within the input sequence, which is necessary because self-attention is inherently permutation-invariant and would otherwise treat a shuffled sequence identically to the original ordering. The sinusoidal functions are chosen for three specific properties identified in the original transformer work [8]. First, each position $t$ receives a unique encoding vector, allowing the model to distinguish any two time steps. Second, the relative offset between two positions $t$ and $t+\Delta$ can be expressed as a linear transformation of the encoding at $t$, which enables the attention heads to learn relative temporal relationships (such as “five trading days ago”) rather than only absolute positions. Third, the deterministic formulation generalizes to sequence lengths not encountered during training without requiring additional learned parameters. The volatile price and volume data are encoded separately in the 17-dimensional feature vector $\mathbf{x}_{i,t}$, which captures the actual market dynamics; the temporal encoding is simply added to this representation so that the model knows the temporal ordering of those observations. Lower encoding dimensions oscillate at higher frequencies, capturing fine-grained positional distinctions between adjacent trading days, while higher dimensions vary more slowly and provide coarser temporal context across the full 252-day input window. The ablation study (Section V-E) confirms that removing temporal encoding degrades MAPE by 18.8%, indicating that positional ordering carries substantial predictive information even though the encoding itself bears no resemblance to price dynamics.
@@ IV-D2 Fine-Tuning Process
where $\alpha_{t}$ is class weight (inverse frequency), $p_{t}$ is predicted probability for correct class, and $\gamma=2$ focuses training on hard examples by down-weighting easily classified instances.
The classification head that maps BERT’s output to a sentiment score consists of three stages. The 768-dimensional [CLS] token representation is first passed through a dropout layer ($p=0.1$) followed by a linear projection to 256 dimensions and a tanh activation, which compresses the representation while preserving the sign structure needed for sentiment polarity. A second dropout layer ($p=0.1$) precedes the final linear projection to three logits corresponding to negative, neutral, and positive classes. Softmax converts these logits to class probabilities $p_{\text{neg}},p_{\text{neu}},p_{\text{pos}}]$, and the continuous sentiment score is computed as $s=p_{\text{pos}}-p_{\text{neg}}\in[-1,+1]$. This formulation produces a score that is exactly zero only when the positive and negative probabilities are equal, and it varies smoothly with confidence, providing a richer signal than a hard class assignment. Fig. 4 illustrates the complete sentiment extraction pipeline from raw input to final score.
@@ IV-E1 Multi-Scale Sentiment Features
Social media activity varies substantially across stocks and calendar days, creating sparsity that requires explicit handling. On trading days where a stock receives fewer than 5 posts, the raw daily score is replaced by the 5-day exponential moving average from preceding days, preventing noisy estimates derived from one or two posts from dominating the signal. On days with no posts at all, all three sentiment features ($S^{(1d)}$, $S^{(5d)}$, $S^{(20d)}$) fall back to their respective moving average values from the most recent day with sufficient coverage. Stocks that consistently receive low social media volume, such as industrials (Caterpillar, Boeing) relative to technology names (Apple, Netflix), naturally produce sentiment scores closer to neutral over time, which the adaptive gating mechanism (Section IV-E) learns to down-weight in favor of price-based features. This design ensures that sparse sentiment data degrades gracefully to a price-only prediction rather than introducing noise.
@@ IV-G Training Process
Throughout all three stages, optimization uses Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The learning rate schedule begins with a linear warmup over 4000 steps followed by cosine decay to zero. A batch size of 32 with gradient accumulation over 4 steps yields an effective batch size of 128, balancing memory constraints with gradient stability.
Hyperparameters were selected through a combination of grid search and established conventions from the transformer literature. Architectural parameters (number of layers, attention heads, model dimension) were searched over $L\in\{4,6,8\}$, $H\in\{4,8\}$, and $d\in\{256,512\}$, with the configuration $L=6$, $H=8$, $d=512$ selected based on lowest validation MAPE. The feed-forward dimension was fixed at $d_{\text{ff}}=4d=2048$ following the standard transformer scaling convention [8]. Dropout rate was searched over $\{0.05,0.1,0.2\}$, with $p=0.1$ yielding the best trade-off between underfitting and overfitting on the validation set. The input sequence length of 252 days (one trading year) was chosen to provide the model with a full annual cycle of price dynamics, and was compared against 120-day and 504-day alternatives on validation data. Learning rates for each training stage were determined through preliminary runs, starting from the range $10^{-5},10^{-3}]$ and narrowing based on training stability and convergence speed. The loss function weights ($\lambda_{1}=1.0$, $\lambda_{2}=0.5$, $\lambda_{3}=0.2$, $\lambda_{4}=10^{-4}$) were tuned by first training with MSE loss alone, then incrementally adding each auxiliary term and adjusting its weight until validation directional accuracy and ranking correlation stabilized. The edge weight balancing parameter $\alpha=0.5$ was searched over $\{0.3,0.5,0.7\}$, with equal weighting of sector and correlation information performing best. Table III summarizes the final hyperparameter configuration for both the node transformer and BERT components.
@@ V-B Evaluation Metrics
The prediction target $y_{i,t+h}$ is the closing price $C_{t+h}$ at horizon $h\in\{1,5,20\}$ trading days. Price-level prediction was chosen as the primary target because it provides intuitive interpretation of forecast accuracy and aligns directly with practical trading applications where position sizing depends on expected price levels rather than percentage changes. A known concern with price-level forecasting is that daily stock prices exhibit strong persistence, meaning that even a trivial model predicting tomorrow’s price as today’s can achieve low MAPE. To guard against this artifact, the evaluation framework includes metrics specifically designed to measure predictive skill beyond persistence: Theil’s U benchmarks against the naive random walk, directional accuracy assesses trend detection independent of magnitude, and the Information Coefficient evaluates cross-sectional ranking. Return-based metrics are also reported (Section V-E) to complement the price-level analysis. Together, these metrics assess model performance from complementary angles, covering magnitude accuracy, scale sensitivity, directional correctness, relative forecasting skill, and cross-sectional ranking. Return-based accuracy is also verified to confirm that results are not artifacts of the evaluation target (Section V-E).
@@ V-C Baseline Models
Eight baseline models spanning statistical, classical machine learning, and deep learning paradigms were trained on identical data splits to ensure fair comparison. All hyperparameters were tuned via grid search on the validation set (2011-2016), with final performance reported on the held-out test set (2017-2025). To ensure comparability across deep learning baselines, all models received the same 17-dimensional feature vector and were trained under equivalent early stopping and regularization protocols.
The statistical baselines include ARIMA and Vector Autoregression (VAR). ARIMA orders $(p,d,q)$ were selected via AIC with $p,q\in[0,5]$ and $d\in[0,2]$, fitting a separate model per stock. VAR extends this to the multivariate setting, modeling cross-stock dependencies with lag order selected via the Bayesian Information Criterion (BIC) up to 10 lags. Both statistical models operate on their own optimal lookback horizons as determined by information criteria, which is standard practice since ARIMA and VAR performance degrades when forced to consume long input windows that introduce noise.
Three classical machine learning models were evaluated. Random Forest used 100 trees with maximum depth 10 and the same 17-dimensional feature vector as the proposed model. Support Vector Regression (SVR) employed a Radial Basis Function (RBF) kernel with regularization parameter $C$ and kernel width $\gamma$ selected via grid search. XGBoost used 1000 estimators with early stopping on validation loss to prevent overfitting, and its gradient boosting framework provides a strong tree-based benchmark. All three classical models used a 60-day lookback window, selected via validation performance from candidates of 20, 60, 120, and 252 days.
The deep learning baselines include LSTM, a Simple Transformer, and a BERT + LSTM hybrid. The LSTM used 2 layers with 128 hidden units and a 60-day lookback window ($\approx$0.5M parameters). During preliminary experiments, the LSTM was also evaluated with 120-day and 252-day lookback windows, but performance degraded at longer horizons due to gradient attenuation over extended sequences, consistent with known limitations of recurrent architectures on very long time series [7]. The 60-day configuration yielded the best validation performance and is therefore reported. The Simple Transformer used 4 layers, 8 attention heads, 256-dimensional embeddings, and a 252-day lookback window matching the proposed model ($\approx$4.2M parameters versus $\approx$21.3M for the full proposed architecture including graph and fusion components). This configuration provides a controlled comparison that isolates the contribution of graph structure and sentiment integration while holding the input horizon and attention mechanism constant. The BERT + LSTM hybrid combined BERT-extracted sentiment features with an LSTM price model through simple concatenation ($\approx$1.1M parameters excluding the frozen BERT encoder), serving as an ablation baseline that tests whether the attention-based fusion mechanism in the proposed architecture provides meaningful improvement over naive feature combination.
The ablation study (Table VII) further addresses baseline fairness by providing graph-free variants of the proposed architecture. The “Without Graph Structure” configuration removes edge weights and inter-stock attention while retaining all other components, effectively reducing the model to a standard transformer with sentiment fusion. This variant isolates the contribution of graph-based relational modeling independent of other architectural advantages.